Starting salaries of a sample of five management majors along with their genders are shown below.

Employee

Salary

(in $1,000s)

Gender

1

30

F

2

28

M

3

22

F

4

26

F

5

19

M

a.

What is the point estimate for the starting salaries of all management majors?

b.

Determine the point estimate for the variance of the population.

c.

Determine the point estimate for the proportion of male employees.
Answer #1

S No x 1 2 3 4 5 EX x=1&ln 30 28 22 26 19 125 25.000 (X-X) 5.00 3.00 -3.00 1.00 -6.00 z(X-X)* $ =z(x-x}*/(n-1)= (X-x) 25.000

a)

point estimate for the starting salaries of all management majors =25.0

b)

point estimate for the variance of the population =20.00

c) point estimate for the proportion of male employees =2/5 =0.4

(since out of 5 employees ;2 are male)